Boyle's law for the expansion of gas PV=C.At a certain instant,the pressure is 2000 lb./ft², the volume is 8 ft³, and the volume increasing at the rate of 6 ft³/mm. Find the rate of change of the pressure at this time


1
Expert's answer
2022-04-14T01:52:08-0400

P1V1 = P2V2

2000 * 8 = P2 * (8+6)

P2 = 1143 lb./ft2

Δ = 2000 – 1143 = 857 lb./ft2*mm